OpenStudy (anonymous):

How to find the base when area is 40 and height is 6?

hero (hero):

of what figure...?

OpenStudy (anonymous):

parallelogram

hero (hero):

area of parallelogram = bh In this case, a = 40, h = 6. so... 40 = 6b 40/6 = b 20/3 = b 6 2/3 = b

OpenStudy (anonymous):

so the base is 6.67